a number is 7 less than another, if 4 times the smaaler number plus twice the larger is 46. find the two numbers?
Answer by nerdybill(7384)

a number is 7 less than another, if 4 times the smaaler number plus twice the larger is 46. find the two numbers?
.
Let x = smaller number
then x-7 = larger number (from:"a number is 7 less than another")
.
Finally, from "4 times the smaaler number plus twice the larger is 46" we get our equation:
4x + 2(x-7) = 47
4x + 2x - 14 = 47
6x - 14 = 47
6x = 61
x = 61/6
x = 10 and 1/6 (smaller number)
.
and,
x-7 = 61/6 - 7 = 61/6 - 42/6 = 19/6
x-7 = 3 and 1/6 (larger number)